A jar contains 10 red marbles numbered 1 to 10 and 8 blue marbles numbered 1 to 8. A marble is drawn at random from the jar. Find the probability of the given event.

(a) The marble is red
Your answer is :
10
18​Correct

(b) The marble is odd-numbered
Your answer is :
9
18​Correct

(c) The marble is red or odd-numbered
Your answer is :

(d) The marble is blue or even-numbered
Your answer is :
1
Expert's answer
2020-07-01T19:04:42-0400

Given:

total number of marbles =18

Formula Probability = number of favorable outcome/ total number of outcomes


Solution :


a)P(red)= number of red /total= 10/18=5/9


b) P(odd)= number of odd/ total= 9/18= 1/2


c)P(red or odd)=P(red)+p(odd)-P(odd and red)

="\frac{10}{18}+\frac{9}{18}-\frac{5}{18}"


="\frac{14}{18}=\frac{7}{9}"

d)P(blue or even)= P(blue)+P(even)-P(blue and even)

="\frac{8}{18}+\frac{9}{18}-\frac{4}{18}"


="\frac{13}{18}"